A patient is given sucrose intravenously. Her venous pressure is 18mmHg, and the elevation difference between the intravenous needle and sucrose bottle is 0.80m. If the rate of sucrose flow is to be 2.0mL/min, what should the diameter of the 4.0 cm long needle be? Assume the sucrose has the same viscosity and density as blood


1
Expert's answer
2021-03-31T09:39:25-0400
"\rho gh=p_1+\frac{8Q\eta L}{\pi r^4}\\
\rho gh=p_1+\frac{Q\eta L}{2\pi d^4}\\
\\
(1050)(9.8)(0.8)=2400+\frac{(3.33\cdot10^{-8})(0.004)(0.04)}{2\pi d^4}\\d=0.11\ mm"